Hi im doing substitution. And i need help on this problem.
y= 5x - 7
-3x - 2y =-12

Answer by lynnlo(4176)

ANSWER: your problem -----y=5x-7
-3x-2y=12
y=3,x=2

THIS IS AN EXAMPLE TO HELP:---------x^2+y^2=25---equation 1
y=x+1-------equation 2
x^2+y^2=25----equation 1
x^2+(x+1)^2=25---substitute x+1 for y
2x^2+2x-24=0-------simplify
2(x+4)(x-3)=0-------factor trinomial
x= -4,(or)x=3---zero product property
this is how to solve a linear quadratic system by substitution